OpenStudy (anonymous):

factor x2 - 5x + 25/4

OpenStudy (anonymous):

its a perfect square. (x-5/2)^2

OpenStudy (theviper):

\[\Huge{\implies x^2-\frac{5x}{2}-\frac{5x}{2}+\frac{25}{4}}\]\[\LARGE{\implies x \left( x-\frac{5}{2} \right)}-\frac{5}{2}\left( x-\frac{5}{2} \right)=??\]

OpenStudy (anonymous):

you should have give him some hint, instead Vamgadu, :)) advise only.

OpenStudy (theviper):

@vamgadu do not give full answer or only solution http://openstudy.com/code-of-conduct